# Aligning by Misaligning: Boundary-aware Curriculum Learning for Multimodal Alignment

### v1

@@ 1 Introduction
Cross–modal representation learning has witnessed rapid progress since CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and their successors demonstrated that contrastive pre-training on web-scale image–text pairs is an effective alternative to costly human annotation. Follow-up models such as ALBEF (Li et al., 2021), BLIP/BLIP-2 (Li et al., 2022, 2023), ViLT (Kim et al., 2021) and GRAM (Cicchetti et al., 2024) further increase sample efficiency by injecting token-level objectives or multi-modal experts(Zhang et al., 2025). These models have promoted the development of fields such as medicine(Fang and Liu, 2025; Liu et al., 2025; Tong et al., 2025) and autonomous systems(Yao et al., 2023; Wang et al., 2025; Li et al., 2025; Lu et al., 2025; Xiao et al., 2025a).
Despite these advances, most existing pipelines share the same implicit assumption: two paired modalities are either perfectly aligned (positive) or entirely unrelated (negative), and the learner’s job is merely to shorten or enlarge their distance (Xin et al., 2024; Fu et al., 2024; Tan et al., 2025; Feng et al., 2025). In practical applications(Jiang et al., 2024), however, cross-modal data often carries subtle mismatches: captions that paraphrase only part of an image, audio tracks that overlap but differ in background context, or video–subtitle pairs where a single phrase is out of sync. These ambiguous negatives—“half-true, half-false” mismatches—are abundant on the web yet overlooked by current training regimes (Yang et al., 2024; Yao et al., 2024).Humans, by contrast, naturally learn from nuanced differences, readily using subtle mismatches as informative cues (Osgood, 1949; Lynn and Bassett, 2020; Hu et al., 2025). Enabling models to similarly leverage these ambiguous negatives is thus essential for more human-like multimodal alignment.
@@
Gap in current practice. Existing alignment methods inadequately handle ambiguous negatives. First, mainstream dual-encoder approaches (e.g., CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), MIL-NCE (Miech et al., 2020b)) sample negatives uniformly, treating obvious mismatches and subtly incorrect captions equally. Second, recent token-aware methods (e.g., ALBEF (Li et al., 2021), BLIP (Li et al., 2022)) discard ambiguous captions through filtering or pseudo-labeling, losing valuable instructional signals, as shown in Fig  1. Finally, existing approaches rely exclusively on static datasets and static loss functions, neglecting dynamically generated, structurally plausible yet semantically ambiguous mismatches (Yan et al., 2024; Xiao and Liu, 2025). Consequently, these methods optimize alignment under idealized conditions, overlooking rich supervisory signals inherent in realistic, partially incorrect data.
Our view. We argue that ambiguous negatives are not merely noise, but a rich supervisory signal: distinguishing almost-correct pairs from truly correct ones is precisely what a robust multimodal model must master when data are scarce or noisy. However, directly training on these borderline cases from the beginning leads to unstable optimisation (Dufumier et al., 2024). The key, therefore, is to schedule the exposure of the learner to increasingly confusing negatives while simultaneously revealing where the confusion arises.
This paper introduces BACL: Boundary-Aware Curriculum with Local Attention. BACL augments any dual-encoder or MoE aligner with two lightweight, fully differentiable components: (i) Boundary-aware Negative Sampler (BNS). A policy network learns to rank candidate negatives by their boundary score and, guided by a logistic schedule, gradually shifts training focus from easy to ambiguous cases. (ii) Contrastive Local Attention (CLA). For every positive pair, CLA compares its cross-attention map with that of the hardest negative and imposes a local mismatch loss that amplifies token pairs where they diverge, forcing the encoder to detect fine-grained misalignment. The main contributions are:
1) We identify ambiguous negatives as an under-explored yet ubiquitous phenomenon in web-scale multimodal corpora and highlight their importance for robust alignment.
2) We propose BACL, the first framework that dynamically generates and exploits near-boundary mismatches via a curriculum sampler and a token-level attention loss.
3) We provide a sharp generalisation theory: under mild assumptions BACL enjoys a $\tilde{\mathcal{O}}(1/n)$ fast rate, whereas uniform sampling suffers an unavoidable $\Omega(\rho/\sqrt{n})$ excess risk (Theorems 4.14.2).
4) Extensive experiments on four large-scale datasets show that BACL improves both global retrieval and fine-grained reasoning, achieving new state-of-the-art results on several benchmarks.
@@ 2 Related Works
Early dual-encoder models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and MIL-NCE (Miech et al., 2020a) learn a shared representation space by drawing uniform in-batch negatives, an approach that proves scalable but treats trivial and near-boundary mismatches alike. Subsequent work improves sample efficiency through stronger vision backbones (ALIGN, OpenCLIP), region-level supervision (UNITER (Chen et al., 2020), ViLT (Kim et al., 2021)), or generative pre-text objectives (ALBEF (Li et al., 2021), BLIP/BLIP-2 (Li et al., 2022, 2023)). Yet, these methods either filter out noisy captions (ALBEF, BLIP) or rely on one-shot max-violation mining (VSE++ (Faghri et al., 2018)), leaving ambiguous negatives under-exploited. Very recent systems such as GRAM (Cicchetti et al., 2024) and Emergence (Tjandrasuwita et al., 2025) incorporate mixture-of-experts or multi-modal masking but still assume a static positive/negative split(Xiao et al., 2025b). Our work departs from this line by dynamically scheduling near-boundary negatives and coupling them with a local attention loss, thereby tightening the decision margin without expensive expert routing or additional labels.
@@
Curriculum learning (Bengio et al., 2009) and self-paced learning (Kumar et al., 2010; Xiao, 2024) propose to present samples in an easy-to-hard order to stabilise optimisation. In computer vision, curricula have been applied to class imbalance (Jiang et al., 2015), structured prediction (Pentina et al., 2015), and more recently to vision-language pre-training DCOT (Wang et al., 2023), where difficulty is measured heuristically (e.g. OT distance). Our Boundary-aware Negative Sampler differs in two aspects: (i) difficulty is defined by a learnable boundary score relative to the current decision margin, and (ii) sampling remains differentiable via Gumbel–Softmax, enabling end-to-end optimisation with the underlying encoders. BACL constitutes the first curriculum framework specifically designed to exploit “half-true, half-false” negatives in multimodal alignment.
@@ 5.1 Experimental Setup
Baselines We benchmark the proposed BACL against five families of vision–language alignment methods: (i) Uniform-negative dual encoders (CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021)). (ii) Single-shot hard-negative mining (VSE++ (Faghri et al., 2018), UNITER (Chen et al., 2020), ALBEF (Li et al., 2021)). (iii) Token-level enhanced pre-training (ViLT (Kim et al., 2021), BLIP (Li et al., 2022), BLIP-2 (Li et al., 2023)). (iv) Curriculum or self-paced alignment (DCOT (Wang et al., 2023)). (v) Multimodal or MoE aligners (Emergence (Tjandrasuwita et al., 2025), CoMM (Dufumier et al., 2024), M3-JEPA (Lei et al., 2024), GRAM (Cicchetti et al., 2024), CLAP (Elizalde et al., 2022) for audio, MIL-NCE (Miech et al., 2020a) for video).
@@ 6 Conclusion
We have presented BACL, a boundary-aware curriculum that converts the ubiquitous yet under-utilised ambiguous negatives into a powerful supervisory signal for multimodal alignment. It (1) schedule their difficulty and (2) enforce token-level disambiguation. The result is a tighter decision margin, provably and empirically. Future work will scale the sampler to billion-image corpora(Chen et al., 2025) and extend the idea to language-only instruction tuning.

### v2

@@ 1 Introduction
Cross–modal representation learning has witnessed rapid progress since CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and their successors demonstrated that contrastive pre-training on web-scale image–text pairs is an effective alternative to costly human annotation (Li et al., 2025b). Follow-up models such as ALBEF (Li et al., 2021), BLIP/BLIP-2 (Li et al., 2022, 2023), ViLT (Kim et al., 2021) and GRAM (Cicchetti et al., 2024) further increase sample efficiency by injecting token-level objectives or multi-modal experts(Zhang et al., 2025, 2024). These models have promoted the development of fields such as natural language processing (Lin et al., 2025), medical diagnosis (Fang and Liu, 2025; Liu et al., 2025; Tong et al., 2025; Wang et al., 2025a; Li et al., 2025c), autonomous systems(Yao et al., 2023; Lu et al., 2025b; Xiao et al., 2025a; Lu et al., 2025a; Li et al., 2025a; Zeng et al., 2025a, b), and other application fields (Xu et al., 2025; Jiang et al., 2025; Tao et al., 2023; Liao et al., 2025; Chan et al., 2026; Sun et al., 2025b; Wang et al., 2025b).
Despite these advances, most existing pipelines share the same implicit assumption: two paired modalities are either perfectly aligned (positive) or entirely unrelated (negative), and the learner’s job is merely to shorten or enlarge their distance (Xin et al., 2024; Fu et al., 2024; Tan et al., 2025; Feng et al., 2025). In practical applications(Jiang et al., 2024; Zhang et al., 2023), however, cross-modal data often carries subtle mismatches: captions that paraphrase only part of an image, audio tracks that overlap but differ in background context, or video–subtitle pairs where a single phrase is out of sync.
These ambiguous negatives—“half-true, half-false” mismatches—are abundant on the web yet overlooked by current training regimes (Yang et al., 2024; Yao et al., 2024).Humans, by contrast, naturally learn from nuanced differences, readily using subtle mismatches as informative cues (Osgood, 1949; Lynn and Bassett, 2020; Hu et al., 2025; Sun et al., 2025a; Chen et al., 2025a). Enabling models to similarly leverage these ambiguous negatives is thus essential for more human-like multimodal alignment.
@@
Gap in current practice.
Existing alignment methods inadequately handle ambiguous negatives. First, mainstream dual-encoder approaches (e.g., CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), MIL-NCE (Miech et al., 2020a)) sample negatives uniformly, treating obvious mismatches and subtly incorrect captions equally. Second, recent token-aware methods (e.g., ALBEF (Li et al., 2021), BLIP (Li et al., 2022)) discard ambiguous captions through filtering or pseudo-labeling, losing valuable instructional signals, as shown in Fig  1. Finally, existing approaches rely exclusively on static datasets and static loss functions, neglecting dynamically generated, structurally plausible yet semantically ambiguous mismatches (Yan et al., 2024; Xiao and Liu, 2025). Consequently, these methods optimize alignment under idealized conditions, overlooking rich supervisory signals inherent in realistic, partially incorrect data (Wang and Zhang, 2024; Lu et al., 2025c).
Our view.
We argue that ambiguous negatives are not merely noise, but a rich supervisory signal: distinguishing almost-correct pairs from truly correct ones is precisely what a robust multimodal model must master when data are scarce or noisy. However, directly training on these borderline cases from the beginning leads to unstable optimisation (Dufumier et al., 2024). The key, therefore, is to schedule the exposure of the learner to increasingly confusing negatives while simultaneously revealing where the confusion arises.
Our solution.
This paper introduces BACL: Boundary-Aware Curriculum with Local Attention. BACL augments any dual-encoder or MoE aligner with two lightweight, fully differentiable components: (i) Boundary-aware Negative Sampler (BNS). A policy network learns to rank candidate negatives by their boundary score and, guided by a logistic schedule, gradually shifts training focus from easy to ambiguous cases. (ii) Contrastive Local Attention (CLA). For every positive pair, CLA compares its cross-attention map with that of the hardest negative and imposes a local mismatch loss that amplifies token pairs where they diverge, forcing the encoder to detect fine-grained misalignment.
Contributions.
The main contributions are:
1) We identify ambiguous negatives as an under-explored yet ubiquitous phenomenon in web-scale multimodal corpora and highlight their importance for robust alignment. We propose BACL, the first framework that dynamically generates and exploits near-boundary mismatches via a curriculum sampler and a token-level attention loss.
2) We provide a sharp generalisation theory: under mild assumptions BACL enjoys a $\tilde{\mathcal{O}}(1/n)$ fast rate, whereas uniform sampling suffers an unavoidable $\Omega(\rho/\sqrt{n})$ excess risk (Theorems 4.14.2).
3) Extensive experiments on four large-scale datasets show that BACL improves both global retrieval and fine-grained reasoning, achieving new state-of-the-art results on several benchmarks.
@@ 2 Related Works
Early dual-encoder models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and MIL-NCE (Miech et al., 2020b) learn a shared representation space by drawing uniform in-batch negatives, an approach that proves scalable but treats trivial and near-boundary mismatches alike (Dong et al., 2024). Subsequent work improves sample efficiency through stronger vision backbones (Jia et al., 2021), region-level supervision (Chen et al., 2020), or generative pre-text objectives (ALBEF (Li et al., 2021), BLIP/BLIP-2 (Li et al., 2022, 2023)). Yet, these methods either filter out noisy captions (ALBEF, BLIP) or rely on one-shot max-violation mining (VSE++ (Faghri et al., 2018)), leaving ambiguous negatives under-exploited. Very recent systems such as GRAM (Cicchetti et al., 2024) and Emergence (Tjandrasuwita et al., 2025) incorporate mixture-of-experts or multi-modal masking but still assume a static positive/negative split(Xiao et al., 2025b). Our work departs from this line by dynamically scheduling near-boundary negatives and coupling them with a local attention loss, thereby tightening the decision margin without expensive expert routing or additional labels.
@@
Curriculum learning (Bengio et al., 2009) and self-paced learning (Kumar et al., 2010; Chen et al., 2024) propose to present samples in an easy-to-hard order to stabilise optimisation. In computer vision, curricula have been applied to class imbalance (Jiang et al., 2015), structured prediction (Pentina et al., 2015), and more recently to vision-language pre-training DCOT (Wang et al., 2023), where difficulty is measured heuristically (e.g. OT distance). Our Boundary-aware Negative Sampler differs in two aspects: (i) difficulty is defined by a learnable boundary score relative to the current decision margin, and (ii) sampling remains differentiable via Gumbel–Softmax, enabling end-to-end optimisation with the underlying encoders. BACL constitutes the first curriculum framework specifically designed to exploit “half-true, half-false” negatives in multimodal alignment.
@@ 5.1 Experimental Setup
Baselines We benchmark the proposed BACL against five families of vision–language alignment methods: (i) Uniform-negative dual encoders (CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021)). (ii) Single-shot hard-negative mining (VSE++ (Faghri et al., 2018), UNITER (Chen et al., 2020), ALBEF (Li et al., 2021)). (iii) Token-level enhanced pre-training (ViLT (Kim et al., 2021), BLIP (Li et al., 2022), BLIP-2 (Li et al., 2023)). (iv) Curriculum or self-paced alignment (DCOT (Wang et al., 2023)). (v) Multimodal or MoE aligners (Emergence (Tjandrasuwita et al., 2025), CoMM (Dufumier et al., 2024), M3-JEPA (Lei et al., 2024), GRAM (Cicchetti et al., 2024), CLAP (Elizalde et al., 2022) for audio, MIL-NCE (Miech et al., 2020b) for video).
@@ 6 Conclusion
We have presented BACL, a boundary-aware curriculum that converts the ubiquitous yet under-utilised ambiguous negatives into a powerful supervisory signal for multimodal alignment. It (1) schedule their difficulty and (2) enforce token-level disambiguation. The result is a tighter decision margin, provably and empirically. Future work will scale the sampler to billion-image corpora(Chen et al., 2025b) and extend our BACL to language-only instruction tuning. Additionally, the boundary-aware curriculum may also be promising for pruning large-scale multimodal models (Zhou et al., 2025).